OpenStudy (sydthekid913):

If f(x) = 3x – 2 and g(x) = 6 – 4x, find f(x) + g(x).

OpenStudy (sydthekid913):

Please help! Will medal and fan!!! Thank you.:)

OpenStudy (ipwnbunnies):

Add similar variables. Add the x terms together. Add the constants together.

OpenStudy (sydthekid913):

I tried but i couldnt get it

OpenStudy (ipwnbunnies):

(3x - 2) + (6 - 4x) =

OpenStudy (sydthekid913):

OOOOOO I got it. Can you answer this one instead?

OpenStudy (sydthekid913):

If given f(x) = x + 4 and g(x) = –2x^2 + 3x – 1, what is (f – g)(x)?

OpenStudy (ipwnbunnies):

.-. No, you can figure it out. Simplify it. 3x - 2 + 6 - 4x = ? 3x - 4x - 2 + 6 = ?

OpenStudy (ipwnbunnies):

(f-g)(x) = f(x) - g(x) Subtract g(x) from f(x).

OpenStudy (sydthekid913):

k